# Diverse Text-to-Image Generation via Contrastive Noise Optimization

### v2

@@ 4.2 Gamma effect: regulated attraction for stable image diversification
$\mathcal{L}^{\gamma}_{\text{CNO}}\coloneqq\mathbb{E}_{i\sim\mathcal{B}}\left[-\log\frac{\exp(\text{sim}(\mathbf{z}_{0|T}^{i,\text{opt}},\mathbf{z}_{0|T}^{i,\text{fixed}})/\gamma\tau)}{\sum_{j\in\mathcal{B}}\exp(\text{sim}(\mathbf{z}_{0|T}^{i,\text{opt}},\mathbf{z}_{0|T}^{j,\text{opt}})/\tau)}\right].$ (7)
@@
$\text{exp}(1/\gamma\tau)=(B-1)\text{exp}(1/\tau).$ (8)
@@ 4.3 Theoretical intuitions
where $I(X;Y)$ is the mutual information between random variables $X$ and $Y$. This inequality implies that minimizing $\mathcal{L}_{\text{InfoNCE}}$ indirectly maximizes $I(Z;Z_{\text{pos}})$, encouraging the learned embedding space to cluster positive pairs. However, this classical relationship does not clarify how negative pairs shape the embedding space—an aspect that is critical in our framework, where negative samples drive diversity.
@@
$\mathcal{L}_{\text{InfoNCE}}\;\geq\;-I(Z;Z_{\mathrm{pos}})+I(Z;Z_{\mathrm{neg}})+\log(B-1),$ (11)
@@
$\mathcal{L^{\gamma}_{\text{InfoNCE}}}\geq-{\frac{1}{\gamma}}I(Z;Z_{\text{pos}})+I(Z;Z_{\text{neg}})+\log(B-1).$ (12)
@@ 5 Experiments
Implementation details. Our experiments are conducted on three distinct pre-trained text-to-image diffusion frameworks: Stable Diffusion v1.5 (SD1.5), SDXL, and SD3. We compare our method with state-of-the-art zero-shot diversity samplers, including Condition-Annealed Diffusion Sampler (CADS) (Sadat et al., 2024) and Particle Guidance (PG) (Corso et al., 2024), as well as the prompt-optimization-based diversity method of Um & Ye (2025b), referred to as DiversityPrompt. All evaluations use text prompts randomly sampled from the MS-COCO (Lin et al., 2014) validation set. For each prompt, we generate 3–5 images, yielding a total of roughly 6–10 K samples.
@@ 5.1 Results
Table 1: Quantitative results of zero-shot diverse samplers. Our proposed method is benchmarked against the standard DDIM sampler and state-of-the-art diversity-enhancing techniques: PG (Corso et al., 2024) and CADS (Sadat et al., 2024). DiversityPrompt refers to the prompt-optimization-based diversity approach developed in Um & Ye (2025b). The evaluation demonstrates that our approach consistently achieves superior performance in diversity metrics, including MSS(↓) and Vendi Score(↑), across Stable Diffusion 1.5, XL, and 3. Notably, it enhances diversity while effectively preserving image quality and prompt fidelity, successfully navigating the fidelity-diversity trade-off by optimizing the initial latent space.
@@
Figure 3 illustrates the quality-diversity trade-off by plotting the Pareto frontiers for our method and key baselines. The plots reveal a clear and compelling advantage for our approach, which establishes a dominant frontier across metrics trained on large-scale human preferences. This is most evident in the PickScore and Image-Reward charts, where our method is strictly superior to competitors like CADS, indicating our outputs are more aesthetically pleasing for any given level of diversity.
While CADS may achieve a marginally higher peak CLIPScore, this metric is known to favor rigid semantic alignment rather than creative or aesthetically superior interpretations. In contrast, our method’s dominance across both PickScore and Image-Reward demonstrates a more intelligent trade-off. It prioritizes what a human user would find visually appealing and contextually appropriate over a mechanical, word-for-word adherence to the prompt. This quantitative strength is mirrored in our qualitative results (see Appendix C.4), where our model uniquely succeeds on complex compositional prompts that cause competitors to fail.
Effect of window size $w$. We conduct an ablation study to investigate the effect of the downsampling window size, $w$, applied to the Tweedie latent shape $\mathbf{z}^{i}_{0|T}$ (Algorithm 1, Line 10). This step is crucial for capturing the global structure of the initial noise prediction while reducing computational cost. We experiment with $w\in\{4,8,16\}$ and compare these against the baseline of $w=64$, which effectively uses the full-resolution latent shape.
The results are illustrated in the PickScore-Vendi Score Pareto frontier in Figure 4. As shown, an aggressive downsampling with $w=4$ leads to a noticeable performance degradation, failing to match the frontier established by larger window sizes. In contrast, moderate downsampling with $w=8$ and $w=16$ achieves highly competitive performance compared to the $w=64$ baseline.
This suggests that moderate downsampling successfully preserves the essential structural information for diversification while benefiting from increased computational efficiency. Excessive downsampling ($w=4$), however, appears to discard critical details necessary for the optimization process. Based on these findings, we select $w=16$ for our main experiments, as it provides the best trade-off between performance and efficiency.
@@ A.1 Proof of Proposition 1
$\mathcal{L}_{\text{InfoNCE}}\;\geq\;-I(Z;Z_{\mathrm{pos}})+I(Z;Z_{\mathrm{neg}})+\log(B-1),$ (13)
@@
In general, we can formulate the infoNCE loss as Eq. (14):
$\mathcal{L}_{\text{InfoNCE}}=-\mathbb{E}_{p(\mathbf{z},\mathbf{z}_{\text{pos}},\mathbf{z}_{\text{neg}})}\left[\log\frac{f(\mathbf{z},\mathbf{z}_{\text{pos}})}{f(\mathbf{z},\mathbf{z}_{\text{pos}})+\sum^{B-1}_{i=1}f(\mathbf{z},\mathbf{z}^{(i)}_{\text{neg}})}\right$]. (14)
@@
$\mathcal{L}_{\text{InfoNCE}}=-\mathbb{E}_{p(\mathbf{z},\mathbf{z}_{p})}\left[\log f(\mathbf{z},\mathbf{z}_{p})\right]+\mathbb{E}_{p(\mathbf{z},\mathbf{z}_{p},\mathbf{z}_{n})}\left[\log\left\{f(\mathbf{z},\mathbf{z}_{p})+\sum^{B-1}_{i=1}f(\mathbf{z},\mathbf{z}^{(i)}_{n})\right\}\right$]. (15)
We can easily assume that $f(\mathbf{z},\mathbf{z}_{p})\geq 0$ by leveraging an exponential function in convention. Then, Eq. (15)  has a lower bound by omitting the positive pair similarity from second term:
$\mathcal{L}_{\text{InfoNCE}}\geq-\mathbb{E}_{p(\mathbf{z},\mathbf{z}_{p})}\left[\log f(\mathbf{z},\mathbf{z}_{p})\right]+\mathbb{E}_{p(\mathbf{z},\mathbf{z}_{n})}\left[\log\sum^{B-1}_{i=1}f(\mathbf{z},\mathbf{z}^{(i)}_{n})\right$]. (16)
According to van den Oord et al. (2019), function $f(\mathbf{z},\mathbf{z^{\prime}})$ estimates the probability density ratio $\frac{p(\mathbf{z}~|~\mathbf{z^{\prime}})}{p(\mathbf{z})}$ related to mutual information maximization. This formulation of $f$ induces the following Eq. (17):
@@ A.2 Proof of Proposition 2
$\mathcal{L^{\gamma}_{\text{InfoNCE}}}\geq-{\frac{1}{\gamma}}I(Z;Z_{\text{pos}})+I(Z;Z_{\text{neg}})+log(B-1).$ (21)
@@
Compared to Equation Eq. (14), similarity function $f(z,z^{\prime})$ is replaced with $f_{\gamma}(z,z^{\prime})=\text{exp}(\frac{sim(z,z^{\prime})}{\gamma})=\{f(z,z^{\prime})\}^{\frac{1}{\gamma}}$. Therefore, Equation Eq. (16) can be rewritten as:
$\mathcal{L}^{\gamma}_{\text{InfoNCE}}\geq-\mathbb{E}_{p(\mathbf{z},\mathbf{z}_{p})}\left[\log\{f(\mathbf{z},\mathbf{z}_{p})\}^{1\over\gamma}\right]+\mathbb{E}_{p(\mathbf{z},\mathbf{z}_{n})}\left[\log\sum^{B-1}_{i=1}f(\mathbf{z},\mathbf{z}^{(i)}_{n})\right$] (22)
$=-{1\over\gamma}\mathbb{E}_{p(\mathbf{z},\mathbf{z}_{p})}\left[\log f(\mathbf{z},\mathbf{z}_{p})\right]+\mathbb{E}_{p(\mathbf{z},\mathbf{z}_{n})}\left[\log\sum^{B-1}_{i=1}f(\mathbf{z},\mathbf{z}^{(i)}_{n})\right$]. (23)
Following similar derivations with Appendix A.1, we can simply show that $\mathcal{L}^{\gamma}_{\text{InfoNCE}}\geq-\frac{1}{\gamma}I(Z;Z_{p})+I(Z;Z_{n})+\log(B-1)$. ∎
@@ B.2 Evaluation metrics
  • Image Quality and Prompt Alignment: To measure the quality and textual relevance of the generated images, we employ a suite of widely-recognized automated metrics.
    • CLIPScore: This metric evaluates the semantic consistency between a generated image and its corresponding text prompt by calculating the cosine similarity of their embeddings from a pre-trained CLIP model (Hessel et al., 2021).
    • PickScore: We use PickScore (Kirstain et al., 2023), a reward model trained on large-scale human preferences, to assess the overall aesthetic quality and prompt alignment of the images.
    • Image-Reward: As a complementary metric, Image-Reward (Xu et al., 2023) is another human-preference-based reward model that provides scores reflecting the general quality of the generated content.
  • Diversity: To evaluate the intra-prompt diversity of the generated images, we utilize two distinct metrics that capture different aspects of variation.
    • Vendi Score: The Vendi Score (Friedman & Dieng, 2023) measures the diversity of a set of samples by analyzing the eigenvalue distribution of their similarity matrix. It provides a holistic assessment of both the variety and balance of the generated images.
    • Mean Pairwise Similarity (MSS): This metric directly quantifies the average similarity between all unique pairs of images generated for a single prompt. We first extract image features using the self-supervised descriptor for image copy detection (SSCD) model (Pizzi et al., 2022). Then, we compute the pairwise cosine similarity matrix of these features and calculate the mean of its off-diagonal elements. A lower MSS value indicates higher diversity, as images in the set are, on average, less similar to one another.
@@ B.3 Hyperparameter settings
Appendix C Additional Analyses and Results

### v3

@@ 4.2 Gamma Effect: Regulated Attraction for Stable Image Diversification
$\mathcal{L}^{\gamma}_{\text{CNO}}\coloneqq\frac{1}{B}\sum_{i=1}^{B}\left[-\log\left(\frac{\exp(f(\hat{\mathbf{z}}_{0|T}^{i},\hat{\mathbf{z}}_{0|T}^{i,\text{ref}})/{(\gamma\tau)})}{\sum_{j=1}^{B}\exp(f(\hat{\mathbf{z}}_{0|T}^{i},\hat{\mathbf{z}}_{0|T}^{j})/\tau)}\right)\right].$ (7)
@@
$\text{exp}(1/{{(\gamma\tau)}})=(B-1)\text{exp}(1/\tau).$ (8)
@@ 4.3 Theoretical Intuitions
where $I(X;Y)$ is the mutual information between random variables $X$ and $Y$. This inequality implies that minimizing $\mathcal{L}_{\text{InfoNCE}}$ indirectly maximizes $I(Z;Z_{\text{pos}})$, encouraging the learned embedding space to cluster positive pairs. However, this classical relationship does not clarify how negative pairs shape the embedding space – an aspect that is critical in our framework, where negative samples drive diversity.
@@
$\mathcal{L}_{\mathrm{InfoNCE}}\;\geq\;-I(Z;Z_{\mathrm{pos}})+I(Z;Z_{\mathrm{neg}})+\log(B-1),$ (11)
@@
${\mathcal{L}^{\gamma}_{\mathrm{CNO}}}\geq-{\frac{1}{\gamma}}I(Z;Z_{\mathrm{pos}})+I(Z;Z_{\mathrm{neg}})+\log(B-1).$ (12)
@@ 5 Experiments
Implementation Details. Our experiments are conducted on three distinct pre-trained text-to-image diffusion frameworks: Stable Diffusion v1.5 (SD1.5), SDXL, and SD3. We compare our method with state-of-the-art zero-shot diversity samplers, including Condition-Annealed Diffusion Sampler (CADS) (Sadat et al., 2024) and Particle Guidance (PG) (Corso et al., 2024), as well as the prompt-optimization-based diversity method of Um and Ye (2025b), referred to as DiversityPrompt. All evaluations use text prompts randomly sampled from the MS-COCO (Lin et al., 2014) validation set. For each prompt, we generate 3–5 images, yielding a total of roughly 6–10 K samples.
@@ 5.1 Results
Table 1: Quantitative results of zero-shot diverse samplers. Our proposed method is benchmarked against standard samplers (DDIM and Flow-Matching ODE) and state-of-the-art diversity-enhancing techniques: PG (Corso et al., 2024) and CADS (Sadat et al., 2024). DiversityPrompt refers to the prompt-optimization-based diversity approach developed in Um and Ye (2025b). The evaluation demonstrates that our approach consistently achieves superior performance in diversity metrics, including MSS(↓) and Vendi Score(↑), across Stable Diffusion 1.5, XL, and 3. Notably, it enhances diversity while effectively preserving image quality and prompt fidelity, successfully navigating the fidelity-diversity trade-off by optimizing the initial latent space.
@@
Figure 4 illustrates the quality-diversity trade-off by plotting the Pareto frontiers for our method and key baselines. The plots reveal a clear and compelling advantage for our approach, which establishes a dominant frontier across metrics trained on large-scale human preferences. This is most evident in the PickScore and Image-Reward charts, where our method is strictly superior to competitors like CADS, indicating our outputs are more aesthetically pleasing for any given level of diversity. While CADS may achieve a marginally higher peak CLIPScore, this metric is known to favor rigid semantic alignment rather than creative or aesthetically superior interpretations. In contrast, our method’s dominance across both PickScore and Image-Reward demonstrates a more intelligent trade-off. It prioritizes what a human user would find visually appealing and contextually appropriate over a mechanical, word-for-word adherence to the prompt. This quantitative strength is mirrored in our qualitative results (see Appendix D.4), where our model uniquely succeeds on complex compositional prompts that cause competitors to fail.
Effect of Window Size $w$. We conduct an ablation study to investigate the effect of the downsampling window size, $w$, applied to the Tweedie latent shape $\hat{\mathbf{z}}^{i}_{0|T}$(Algorithm 1, Line 10). This step is crucial for capturing the global structure of the initial noise prediction while reducing computational cost. We experiment with $w\in\{4,8,16\}$ and compare these against the baseline of $w=64$, which effectively uses the full-resolution latent shape.
The results are illustrated in the PickScore-Vendi Score Pareto frontier in Figure 3. As shown, an aggressive downsampling with $w=4$ leads to a noticeable performance degradation, failing to match the frontier established by larger window sizes. In contrast, moderate downsampling with $w=8$ and $w=16$ achieves highly competitive performance compared to the $w=64$ baseline. This suggests that moderate downsampling successfully preserves the essential structural information for diversification while benefiting from increased computational efficiency. Excessive downsampling ($w=4$), however, appears to discard critical details necessary for the optimization process. Based on these findings, we select $w=16$ for our main experiments, as it provides the best trade-off between performance and efficiency.
@@ 6 Conclusion
Acknowledgments
This research was supported by the National Research Foundation of Korea (NRF) under Grant RS-2024-00336454, and the AI Computing Infrastructure Enhancement (GPU Rental Support) User Support Program funded by the Ministry of Science and ICT (MSIT), Republic of Korea (RQT-25-120217). This work was also supported by the Institute of Information & Communications Technology Planning & Evaluation (IITP) grant funded by the Korean government (MSIT) (No. RS-2019-II190075, Artificial Intelligence Graduate School Program(KAIST); No. RS-2025-02304967, AI Star Fellowship (KAIST); No. RS-2025-02219317, AI Star Fellowship (Kookmin University)).
@@ A.1 Proof of Proposition 1
$\mathcal{L}_{\mathrm{InfoNCE}}\;\geq\;-I(Z;Z_{\mathrm{pos}})+I(Z;Z_{\mathrm{neg}})+\log(B-1),$ (13)
@@
In general, we can formulate the infoNCE loss as Eq. (14) by setting $g(\cdot,\cdot)=\exp(f(\cdot,\cdot)/\tau)$ from Eq. (5):
$\mathcal{L}_{\text{InfoNCE}}=-\mathbb{E}_{p(\mathbf{z},\mathbf{z}_{\text{pos}},\mathbf{z}_{\text{neg}})}\left[\log\frac{{g}(\mathbf{z},\mathbf{z}_{\text{pos}})}{{g}(\mathbf{z},\mathbf{z}_{\text{pos}})+\sum^{B-1}_{i=1}{g}(\mathbf{z},\mathbf{z}^{(i)}_{\text{neg}})}\right$]. (14)
@@
$\mathcal{L}_{\text{InfoNCE}}=-\mathbb{E}_{p(\mathbf{z},\mathbf{z}_{p})}\left[\log{g}(\mathbf{z},\mathbf{z}_{p})\right]+\mathbb{E}_{p(\mathbf{z},\mathbf{z}_{p},\mathbf{z}_{n})}\left[\log\left\{{g}(\mathbf{z},\mathbf{z}_{p})+\sum^{B-1}_{i=1}{g}(\mathbf{z},\mathbf{z}^{(i)}_{n})\right\}\right$]. (15)
Given that $g(\cdot,\cdot)$ is non-negative due to its exponential form, Eq. (15)  has a lower bound by omitting the positive pair similarity from second term:
$\mathcal{L}_{\text{InfoNCE}}\geq-\mathbb{E}_{p(\mathbf{z},\mathbf{z}_{p})}\left[\log{g}(\mathbf{z},\mathbf{z}_{p})\right]+\mathbb{E}_{p(\mathbf{z},\mathbf{z}_{n})}\left[\log\sum^{B-1}_{i=1}{g}(\mathbf{z},\mathbf{z}^{(i)}_{n})\right$]. (16)
According to van den Oord et al. (2019), function ${g}(\mathbf{z},\mathbf{z^{\prime}})$ estimates the probability density ratio $\frac{p(\mathbf{z}~|~\mathbf{z^{\prime}})}{p(\mathbf{z})}$ related to mutual information maximization. This formulation of g{g} induces the following Eq. (17):
@@ A.2 Proof of Proposition 2
$\mathcal{L^{\gamma}_{\mathrm{CNO}}}\geq-{\frac{1}{\gamma}}I(Z;Z_{\mathrm{pos}})+I(Z;Z_{\mathrm{neg}})+log(B-1).$ (21)
@@
Compared to Equation Eq. (14), similarity function ${g}(z,z^{\prime})$ is replaced with ${g}_{\gamma}(z,z^{\prime})=\text{exp}(\frac{f(z,z^{\prime})}{\gamma{\tau}})=\{{g}(z,z^{\prime})\}^{\frac{1}{\gamma}}$. Therefore, Equation Eq. (16) can be rewritten as:
$\mathcal{L}^{\gamma}_{\text{CNO}}\geq-\mathbb{E}_{p(\mathbf{z},\mathbf{z}_{p})}\left[\log\{{g}(\mathbf{z},\mathbf{z}_{p})\}^{1\over\gamma}\right]+\mathbb{E}_{p(\mathbf{z},\mathbf{z}_{n})}\left[\log\sum^{B-1}_{i=1}{g}(\mathbf{z},\mathbf{z}^{(i)}_{n})\right$] (22)
$=-{1\over\gamma}\mathbb{E}_{p(\mathbf{z},\mathbf{z}_{p})}\left[\log{g}(\mathbf{z},\mathbf{z}_{p})\right]+\mathbb{E}_{p(\mathbf{z},\mathbf{z}_{n})}\left[\log\sum^{B-1}_{i=1}{g}(\mathbf{z},\mathbf{z}^{(i)}_{n})\right$]. (23)
Following similar derivations with Appendix A.1, we can simply show that $\mathcal{L}^{\gamma}_{\text{CNO}}\geq-\frac{1}{\gamma}I(Z;Z_{p})+I(Z;Z_{n})+\log(B-1)$. ∎
@@ B.2 Evaluation Metrics
  • Image Quality and Prompt Alignment. To measure the quality and textual relevance of the generated images, we employ a suite of widely-recognized automated metrics.
    • CLIPScore. This metric evaluates the semantic consistency between a generated image and its corresponding text prompt by calculating the cosine similarity of their embeddings from a pre-trained CLIP model (Hessel et al., 2021).
    • PickScore. We use PickScore (Kirstain et al., 2023), a reward model trained on large-scale human preferences, to assess the overall aesthetic quality and prompt alignment of the images.
    • Image-Reward. As a complementary metric, Image-Reward (Xu et al., 2023) is another human-preference-based reward model that provides scores reflecting the general quality of the generated content.
  • Diversity. To evaluate the intra-prompt diversity of the generated images, we utilize two distinct metrics that capture different aspects of variation.
    • Vendi Score. The Vendi Score (Friedman and Dieng, 2023) measures the diversity of a set of samples by analyzing the eigenvalue distribution of their similarity matrix. It provides a holistic assessment of both the variety and balance of the generated images.
    • Mean Pairwise Similarity (MSS). This metric directly quantifies the average similarity between all unique pairs of images generated for a single prompt. We first extract image features using the self-supervised descriptor for image copy detection (SSCD) model (Pizzi et al., 2022). Then, we compute the pairwise cosine similarity matrix of these features and calculate the mean of its off-diagonal elements. A lower MSS value indicates higher diversity, as images in the set are, on average, less similar to one another.
@@ B.3 Hyperparameter Settings
Appendix C Further Analyses and Discussions
C.1 Stepwise Mechanism of Constrastive Noise Optimization
To clarify the exact role of the attraction coefficient $\gamma$ and the gradient dynamics within our framework, we provide a stepwise breakdown using a minimal batch example.
Loss Mechanism: Attraction and Repulsion. Consider a minimal batch of size $B=2$. Let $\mathcal{L}_{1}$ denote the loss for the first initial noise vector $\mathbf{z}_{T}^{1}$. Given a fixed reference anchor $\hat{\mathbf{z}}_{0|T}^{1,\text{ref}}$ and the similarity function $\text{sim}(\cdot,\cdot)$, the loss decomposes into two distinct forces:
$\mathcal{L}_{1}$ $=-\log\frac{\exp(\text{sim}(\hat{\mathbf{z}}_{0|T}^{1},\hat{\mathbf{z}}_{0|T}^{1,\text{ref}})/(\gamma\tau))}{\sum_{j=1}^{2}\exp(\text{sim}(\hat{\mathbf{z}}_{0|T}^{1},\hat{\mathbf{z}}_{0|T}^{j})/\tau)}$
$=\underbrace{-\frac{\text{sim}(\hat{\mathbf{z}}_{0|T}^{1},\hat{\mathbf{z}}_{0|T}^{1,\text{ref}})}{\gamma\tau}}_{\text{(A) Attraction}}+\underbrace{\log\left(\exp(\text{sim}(\hat{\mathbf{z}}_{0|T}^{1},\hat{\mathbf{z}}_{0|T}^{1})/\tau)+\exp(\text{sim}(\hat{\mathbf{z}}_{0|T}^{1},\hat{\mathbf{z}}_{0|T}^{2})/\tau)\right)}_{\text{(B) Repulsion}}$
(A) Attraction. This term encourages alignment between the current estimate $\hat{\mathbf{z}}_{0|T}^{1}$ and its fixed reference $\hat{\mathbf{z}}_{0|T}^{1,\text{ref}}$, preserving semantic fidelity to the original concept.
(B) Repulsion. It pushes $\hat{\mathbf{z}}_{0|T}^{1}$ away from other samples in the batch (e.g., $\hat{\mathbf{z}}_{0|T}^{2}$), enforcing diversity by maximizing semantic distance.
Role of Gamma ($\gamma$). The coefficient $\gamma$ serves as a regulator for the fidelity-diversity trade-off by exclusively scaling the attraction term. As derived in Proposition 2, the effective objective is lower-bounded by:
$\mathcal{L}^{\gamma}_{\text{CNO}}\geq-\frac{1}{\gamma}I(Z;Z_{\text{pos}})+I(Z;Z_{\text{neg}})+\log(B-1).$
This reveals that $\gamma$ modulates the strength of the positive mutual information. Specifically:
  • Decreasing $\gamma<1$: Amplifies the attraction force ($1/\gamma>1$). This is crucial for larger batch sizes (e.g., $B=3$), where the cumulative repulsion from B1B-1 negative pairs can overshadow the single attraction term. A lower $\gamma$ restores balance, preventing the sample from drifting too far from the anchor.
  • Increasing $\gamma>1$: Dampens the attraction, allowing the repulsion term to dominate. This promotes greater diversity but risks reducing fidelity.
Practically, we find that setting $\gamma\approx(\tau\ln(B-1)+1)^{-1}$ provides a robust baseline for balancing these forces across varying batch sizes.
Gradient Flow and the Stop-Gradient Strategy. To optimize $\mathbf{z}_{T}$ efficiently, we analyze the gradient flow. For the loss $\mathcal{L}_{1}$ defined above:
  • Active Gradient Flow: Gradients propagate through $\hat{\mathbf{z}}_{0|T}^{1}$ in both numerator and denominator, driving it toward the anchor and away from negatives.
  • No Flow to Anchor: The reference $\hat{\mathbf{z}}_{0|T}^{1,\text{ref}}$ is fixed; thus, no gradients flow through this term, ensuring it remains a stable guidepost.
Crucially, to backpropagate from $\hat{\mathbf{z}}_{0|T}$ to $\mathbf{z}_{T}$, we utilize a stopgrad operation on the diffusion model output $\bm{\epsilon}_{\theta}$. By the chain rule:
$\nabla_{\mathbf{z}_{T}}\mathcal{L}$ $=\left(\frac{\partial\hat{\mathbf{z}}_{0|T}}{\partial\mathbf{z}_{T}}\right)^{T}\nabla_{\hat{\mathbf{z}}_{0|T}}\mathcal{L}$
$=\frac{1}{\sqrt{\bar{\alpha}_{T}}}\left(\mathbf{I}-\sqrt{1-\bar{\alpha}_{T}}\frac{\partial\bm{\epsilon}_{\theta}(\mathbf{z}_{T})}{\partial\mathbf{z}_{T}}\right)^{T}\nabla_{\hat{\mathbf{z}}_{0|T}}\mathcal{L}.$
Calculating the full Jacobian $\partial\bm{\epsilon}_{\theta}/\partial\mathbf{z}_{T}$ is computationally prohibitive. By applying stopgrad, we set this term to zero, simplifying the update to:
$\nabla_{\mathbf{z}_{T}}\mathcal{L}\approx\frac{1}{\sqrt{\bar{\alpha}_{T}}}\cdot\mathbf{I}\cdot\nabla_{\hat{\mathbf{z}}_{0|T}}\mathcal{L}.$
This approximation effectively updates the noise directly in the direction of the semantic gradient. This strategy mirrors the effective Jacobian approximation used in Score Distillation Sampling (SDS) (Poole et al., 2023) and NoiseRefine (Ahn et al., 2024), ensuring stable and efficient guidance without the cost of backpropagating through the U-Net.